### v1

# Degeneracy Enhancement of Neutron-Antineutron Oscillation in Neutron Star

Abstract
We explore the fermion oscillation in a degenerate environment. The direct consequence is introducing a Pauli blocking factor $1-f_{i}$, where $f_{i}$ is the phase space distribution function, for each intermediate mass eigenstate during propagation. It is then much easier for a state with larger existing fraction or density to oscillate into other states with less degeneracy while the reversed process is not enhanced. This can significantly modify the oscillation behaviors. We apply this degenerate fermion oscillation to a concrete scenario of neutron-antineutron oscillation in neutron star. It turns out antineutrons receive a standing fraction to annihilate with the environmental neutrons. The subsequent neutron star heating can put an extremely stringent bound on the baryon number violating cross mass term between neutron and antineutron.
Introduction – The Grand Unification Theory (GUT) Georgi:1974sy ; Pati:1974yy ; Mohapatra:1974hk ; Fritzsch:1974nn is a fundamental theory that seeks to unify the electromagnetic, weak, and strong forces Croon:2019kpe . In addition to proton decay, the neutron-antineutron ($n\bar{n}$) oscillation is also an important prediction of some GUT models Mohapatra:1996pu . While proton decay violates the baryon number by one, $n\bar{n}$ oscillation appears to be a higher-order process with baryon number violation by two. These could be an important source for the baryon asymmetry observed in our Universe Fukugita:2002hu ; Bodeker:2020ghk .
Previous experiments on $n\bar{n}$ oscillation have primarily focused on the processes happening inside atomic nuclei Kamiokande:1986pyk ; Chung:2002fx ; Friedman:2008es or in a vacuum environment Baldo-Ceolin:1994hzw . These experiments have provided valuable constraints on the oscillation parameters, in particular a lower limit on the oscillation time scale $\tau_{n\bar{n}}\sim 1/\delta m$, where $\delta m$ is the off-diagonal mass matrix element that stands for neutron-antineutron mixing, up to the level of $10^{8}$ seconds and hence a lower limit $10^{4.5}$ GeV for the GUT scale Buccella:1999wy ; Babu:2012vc . In comparison, the proton decay sensitivity can reach $10^{34}$ years for lifetime and $10^{14}$ GeV for the GUT scale Dev:2022jbf ; Ohlsson:2023ddi .
The neutron star (NS) has been used as a natural laboratory for studying exotic new physics phenomena Raffelt:1996wa . It is also possible to use NS for probing the $n\bar{n}$ oscillation Krishan:ApSS82 ; Buccella:1987pv . However, the previous studies show that the $n\bar{n}$ oscillation would not significantly impact the NS evolution Buccella:1987pv ; Buccella:1999wy . In comparison with the terrestrial experiments that can detect individual annihilation signal induced by the $n\bar{n}$ oscillation, the NS heating can only be sensitive to the macroscopic effect. The NS constraint ceased from study and was only briefly commented by McKeen:2021jbh and almost touched by Berezhiani:2020vbe .
In this paper, we first point out the degenerate scenario of fermion oscillation in a general context and then apply it to the $n\bar{n}$ oscillation inside NS. Since the neutron matter inside NS is highly degenerate, it is much easier for neutron to oscillate into antineutron but the reverse process is much more difficult. This enhancement of $n\bar{n}$ oscillation is very significant and allows NS to provide much stronger constraint than previously expected.
Fermion Oscillation with Degeneracy – The formalism for fermion oscillation in vacuum is well established for neutrinos Pontecorvo:1957cp ; MNS ; P67 . The initial flavor state $|\alpha\rangle$ can be decomposed into a linear combination of mass eigenstates $|i\rangle$, $|\alpha\rangle=\sum_{i}U^{*}_{\alpha i}|i\rangle$ with $U_{\alpha i}$ being the mixing matrix element Kayser:2004wmw . With definite dispersion relation and hence propagating as plane wave, the mass eigenstate evolves independently, $|i(t,L)\rangle=e^{i\phi_{i}(t,L)}|i\rangle$ where $\phi_{i}(t,L)$ is the corresponding evolution phase over time interval $t$ and distance $L$. Correspondingly, the original flavor state $|\alpha(t,L)\rangle$ also evolves with time and distance. For measurement also in a flavor state $|\beta\rangle$, the amplitude $A_{\beta\alpha}\equiv\langle\beta|\alpha(t,L)\rangle$ is no longer simply $\langle\beta|\alpha\rangle=\delta_{\alpha\beta}$ but
$A_{\beta\alpha}=\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}.$ (1)
With different evolution phases $\phi_{i}\neq\phi_{j}$, their interference leads to oscillatory behavior in the transition probability $P_{\alpha\beta}\equiv|A_{\beta\alpha}|^{2}$. We have kept our notations general in order to cover neutrino and other fermion oscillations. For neutrino oscillation, the individual evolution phase is $\phi_{i}=m^{2}_{i}L/2E_{\nu}$ and the oscillation phase is $\Delta\phi_{ij}=\Delta m^{2}_{ij}L/2E_{\nu}$ with $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$. One may see that the neutrino oscillation phenomena is actually a quantum interference over macroscopic distance.
The vacuum fermion oscillation is established by considering fermion propagation on the vacuum state. However, if the fermion propagates through a degenerate environment fully filled with fermions of the same type, the amplitude into a certain mass eigenstate can be suppressed due to the Pauli exclusion principle which will significantly affect the oscillation probabilities. In a degenerate environment, the vacuum state $|0\rangle$ should be replaced by the degenerate environment state $|\Omega\rangle$.
In a concrete process, the fermion in propagation is created by a creation operator $a^{\dagger}_{\bm{p}i}$ and then propagate in the dense environment before being annihilated by $a_{\bm{p}i}$. The corresponding amplitude takes the form as $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle$. For simplicity, we have omitted other parts such as the production and detection vertices. Note that the creation and annihilation operators are always associated with a physical mass eigenstate that has definite dispersion relation. If the mass eigenstate propagate over distance, the evolution phase $e^{i\phi_{i}(t,L)}$ should also be taken into consideration.
The product of creation and annihilation operators is closely related to the particle momentum density, $\hat{n}_{\bm{p}i}\equiv a^{\dagger}_{\bm{p}i}a_{\bm{p}i}$. When applying to the environmental state $|\Omega\rangle$, the number density operator $\hat{n}_{\bm{p}i}$ should produce the particle momentum density, $\hat{n}_{\bm{p}i}|\Omega\rangle=n_{\bm{p}i}|\Omega\rangle$ with $n_{\bm{p}i}\equiv\int\mathrm{d}^{3}\bm{x}f_{i}(\bm{x},\bm{p})$ where $f_{i}(\bm{x},\bm{p})$ is the corresponding phase space distribution function. For a fermion, the creation and annihilation operators anticommute as, $\{a_{\bm{p}_{1}i},a_{\bm{p}_{2}i}^{\dagger}\}=(2\pi)^{3}\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$. Consequently, the product $a_{\bm{p}i}a^{\dagger}_{\bm{p}i}$ gives the Pauli blocking factor,
$\displaystyle{a_{\bm{p}i}a_{\bm{p}i}^{\dagger}|\Omega\rangle=\int\mathrm{d}^{3}\bm{x}[1-f_{i}(\bm{x},\bm{p})]}|\Omega\rangle,$ (2)
where the $\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$ has been removed by momentum integration in the second quantized field operator, $\nu(x)\sim\int\mathrm{d}^{4}p(aue^{-ip\cdot x}+b^{\dagger}ve^{ip\cdot x})$.
Taking the neutrino oscillation as a concrete example, the whole scattering amplitude of production, propagation, and detection takes the form as,
$\mathcal{M}$ $=\langle f,\Omega|(\bar{\ell}_{\beta}\gamma^{\mu}P_{L}W_{\mu}^{-}\nu_{\beta})(\bar{\nu}_{\alpha}\gamma^{\nu}P_{L}\ell_{\alpha}W_{\nu}^{+})|i,\Omega\rangle$ (3a)
$\equiv\mathcal{M}_{d}\left[\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle\right]\mathcal{M}_{p},$ (3b)
where $W^{\pm}$ is the $W$ gauge boson and $\ell_{\alpha}$ the charged lepton of flavor $\alpha$. For convenience, the weak coupling constant has been omitted. The two neutrino spinors $\bar{u}$ and $u$ have been absorpted into the production and detection matrix elements $\mathcal{M}_{p}$ and $\mathcal{M}_{d}$, respectively. In between, the creation and annihilation operators sandwiched by the thermal environment state $|\Omega\rangle$ would give the Pauli exclusion factor, $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle=1-f_{i}$. Although illustrated with neutrino oscillation, this feature should apply generally to any fermion oscillation.
The oscillation amplitude in Eq. (1) then generalizes to
$A_{\beta\alpha}=\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}(1-f_{i}).$ (4)
The Pauli suppression factor $1-f_{i}$ represents the degeneracy effect brought by the environment. With an empty environment as vacuum ($f_{i}=0$), the amplitude reduces to the vacuum oscillation case as shown in Eq. (1). On the other hand, a fully degenerate Fermi gas $f_{i}=1$ does not allow oscillation to occupy the corresponding mass eigenstate.
With Pauli suppression factor, the amplitude in Eq. (4) can no longer reflect unitarity. Sticking to the usual definition of oscillation probability, $P_{\alpha\beta}\equiv|A_{\beta\alpha}|^{2}$, the total probability, $P_{\text{tot},\alpha}\equiv\sum_{\gamma}P_{\alpha\gamma}\neq 1$, is no longer one. It seems that the oscillating fermion would somehow disappear which is contradictory to the intuitive physical picture. However, oscillation in a degenerate environment does not introduce any mechanism for a particle to disappear. So unitarity has to be maintained.
To resolve this contradiction, one needs to resort to the more intuitive definition of oscillation probability. The original definition of probability or chance is actually the fraction of events. To be more exact, the oscillation probability $P_{\alpha\beta}$ should be the ratio between the number of $\beta$ flavor particles that can be detected after oscillation and the original number of $\alpha$ flavor ones without oscillation. It is then necessary to normalize the oscillation probability as $\mathcal{P}_{\alpha\beta}\equiv|A_{\alpha\beta}|^{2}/P_{\text{tot}}$ by hand. Note that the amplitude $A_{\alpha\beta}$ in Eq. (4) does not reduce to $\delta_{\alpha\beta}$ in the zero distance limit since the Pauli blocking is already in effect even if the oscillation does not really happen. In a degenerate environment, an $\alpha$ flavor state from production cannot contain the existing mass eigenstate.
Neutron-Antineutron Oscillation with Degeneracy – Being neutral particles, neutron and antineutron can have cross mass term Phillips:2014fgb . To be exact, the neutron and antineutron Hamiltonian is a $2\times 2$ matrix,
$H\approx\begin{pmatrix}H_{11}&\delta m\\ \delta m&H_{22}\end{pmatrix},$ (5)
in the $(n,\bar{n})$ basis. The diagonal elements $H_{11}$ and $H_{22}$ are the neutron and antineutron energies Phillips:2014fgb . The cross mass $\delta m$ will induce the $n$-$\bar{n}$ mixing, $\tan 2\theta=2\delta m/(H_{22}-H_{11})$,
$U=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix},\quad\begin{pmatrix}n_{1}\\ n_{2}\end{pmatrix}=U\begin{pmatrix}n\\ \bar{n}\end{pmatrix},$ (6)
where $n_{1}$ and $n_{2}$ are the two mass eigen-fields. Typically, the mixing angle $\theta$ is highly suppressed by the tiny $\delta m$. For simplicity, one may denote the sine and cosine of the mixing angle as $(s,c)\equiv(\sin\theta,\cos\theta)$. The energy difference between the mass eigenstates is simply denoted as $\Delta E\equiv\sqrt{(H_{11}-H_{22})^{2}+4\delta m^{2}}$ Phillips:2014fgb . So the $n\bar{n}$ oscillation probabilities are
$P_{n\bar{n}}$ $=c^{2}s^{2}(1-f_{1})^{2}+s^{2}c^{2}(1-f_{2})^{2}$
$-2c^{2}s^{2}(1-f_{1})(1-f_{2})\cos(\Delta Et),$ (7a)
$P_{nn}$ $=c^{4}(1-f_{1})^{2}+s^{4}(1-f_{2})^{2}$
$+2c^{2}s^{2}(1-f_{1})(1-f_{2})\cos(\Delta Et),$ (7b)
where $f_{1}$ and $f_{2}$ are the phase space distribution functions of the two mass eigenstates. Note that the time dependence only appears in the $\cos(\Delta Et)$ term and the norminal total probability $P_{\rm tot}\equiv P_{nn}+P_{n\bar{n}}$ is indeed time-independent which is a necessary feature for our argument of probability normalization/definition above. The two probabilities reduce to the non-degenerate case with both $f_{1}$ and $f_{2}$ approaching 00.
The normalized oscillation probability should take the form as,
$\mathcal{P}_{n\bar{n}}$ $=\frac{c^{2}s^{2}(1-f_{1})^{2}+c^{2}s^{2}(1-f_{2})^{2}}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}}$
$-\frac{2(1-f_{1})(1-f_{2})c^{2}s^{2}\cos(\Delta Et)}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}},$ (8)
with $\mathcal{P}_{nn}\equiv P_{nn}/P_{\rm tot}$ being omitted for simplicity but can be readily reconstructed. The total probability $P_{\rm tot}$ in the denominator reduces to 1 in the non-degenerate limit. It is interesting to see that the phase space distribution function appears in the denominator as $1-f_{i}$. In a dense environment with degeneracy, the phase space distribution function $f_{i}$ can approach 1 below the Fermi surface. It is then possible for the oscillation probabilities to receive enhancement from degeneracy.
With a tiny mixing angle $\theta$ Phillips:2014fgb , $s\ll 1$, the mass eigenstate $n_{1}$ is approximately the neutron state $n$ that dominates the matter inside NS. Consequently, both $1-f_{1}$ and $f_{2}$ are negligibly small. Between the two tiny terms $s$ and $1-f_{1}$ in the denominator, the case $s\gg 1-f_{1}$ would lead to significant transition rate $\mathcal{P}_{n\bar{n}}\approx c^{2}$ which means degenerate neutrons would instantly transit to antineutrons. The situation $s\sim 1-f_{1}$ has also the same conclusion. Since neutron stars are readily observed and hence should be stable, only the remaining case $s\ll 1-f_{1}$ can happen. This already indicates that the degenerate $n\bar{n}$ oscillation in NS can give a very much enhanced constraint.
To be more exact, in the limit of $s\ll 1-f_{1}\ll 1$ and $f_{2}\ll 1$, the transition rate from neutron to antineutron becomes
$\mathcal{P}_{n\bar{n}}\approx\frac{s^{2}}{(1-f_{1})^{2}}\equiv s^{2}_{\rm eff},\qquad\mathcal{P}_{nn}\approx c^{2}.$ (9)
It is interesting to observe that the transition rate $\mathcal{P}_{n\bar{n}}$ is enhanced by degeneracy. This is consistent with the intuitive picture that it is much easier for neutrons in a degenerate neutron environment to transit to other particles without Pauli blocking. The effective mixing $s^{2}_{\rm eff}$ should be much larger than the vacuum counterpart $s^{2}$. Switching neutron with antineutron and hence $f_{1}$ with $f_{2}$, the reversed transition probability $\mathcal{P}_{\bar{n}n}\approx s^{2}$ is not enhanced. The bidirectional oscillation has imbalanced rates to derive the overall transition from the degenerate neutrons to the non-degenerate antineutrons.
Standing Fraction of Antineutron in Neutron Star – Another important feature is that the time dependence of the $n\bar{n}$ oscillation probability $\mathcal{P}_{n\bar{n}}$, which is $2c^{2}s^{2}\sin^{2}(\Delta Et/2)$ for the vacuum case, is now suppressed by extra $1-f_{1}$. In other words, the transition happens instantaneously. In an extremely degenerate environment such as NS, neutrons will transit into antineutrons so rapidly that the ratio between their distribution functions $f_{\bar{n}}/f_{n}$ stays the same,
$R(\bm{p})\equiv\frac{f_{\bar{n}}}{f_{n}}=\frac{\mathcal{P}_{n\bar{n}}}{\mathcal{P}_{nn}}\approx\frac{\tan^{2}\theta}{[1-f_{1}(\bm{p})]^{2}},$ (10)
which actually reduces to the effective mixing $s^{2}_{\rm eff}$. With momentum dependence inherited from $f_{1}(\bm{p})$, $s^{2}_{\rm eff}$ does not have a fixed value. Even directly taking the zero time limit of Eq. (8), $\mathcal{P}_{n\bar{n}}(t=0)\approx s^{2}(f_{1}-f_{2})^{2}/(1-f_{1})^{2}$ where the approximation only applies for the denominator, Eq. (10) can be readily reproduced with $(f_{1}-f_{2})^{2}\approx 1$. In other words, the production process is already affected by the degenerate environment as pointed out earlier. The neutron transition to antineutron can instantly happen and remain the same antienutron fraction over time.
Being in thermal equilibrium as maintained inside NS, neutrons follow the Fermi-Dirac distribution, $f_{1}(\bm{p})=1/[e^{(\varepsilon_{n}(\bm{p})-\mu)/T}+1]$ where $\varepsilon_{n}(\bm{p})=\sqrt{\bm{p}^{2}+m^{2}}$ is the neutron energy with momentum $\bm{p}$ and mass $m$. With a very low internal temperature $T\sim 7$ keV at the age of $10^{5}$ years Potekhin:2020ttj , the corresponding neutron Fermi momentum $k_{F}=(3\pi^{2}n_{n})^{1/3}$ and chemical potential $\mu_{n}=\sqrt{k_{F}^{2}+m^{2}}$ can be readily derived from the neutron density $n_{n}$ assuming full degeneracy. We consider a typical NS with mass $M=1.4M_{\odot}$ (solar mass $M_{\odot}$) and radius $R_{n}=10$ km Lattimer:2000nx . To match the corresponding high density $n_{n}\approx 4\times 10^{38}$ cm-3 Potekhin:2015qsa , which is also assumed to be constant inside NS, the difference of chemical potential $\mu$ from the neutron mass $m$, $\mu-m\approx 101.6$ MeV, is much larger than temperature $T$ Brandes:2023bob . It is fully consistent to say that the distribution function is very close to 1 below the Fermi surface, $1-f_{1}(\bm{p})\approx e^{[\varepsilon_{n}(\bm{p})-\mu]/T}\ll 1$.
Since the first mass eigenstate is mainly neutron, $n_{1}=c\,n+s\,\bar{n}$ according to Eq. (6), they share roughly the same phase space distribution, $f_{n}\approx f_{1}$. The standing fraction of antineutron as shown in Eq. (10) determines the antineutron phase space distribution, $f_{\bar{n}}(\bm{p})=R(\bm{p})f_{n}(\bm{p})\approx s^{2}f_{1}(\bm{p})/[1-f_{1}(\bm{p})]$ and the antineutron number density $n_{\bar{n}}=\int\frac{g_{s}}{(2\pi)^{3}}\mathrm{d}^{3}\bm{p}f_{\bar{n}}(\bm{p})$ can be analytically obtained,
$n_{\bar{n}}=\frac{m^{2}T}{2\pi^{2}}e^{\frac{\mu}{T}}\left[2K_{2}\left(\frac{m}{T}\right)+e^{\frac{\mu}{T}}K_{2}\left(\frac{2m}{T}\right)\right]s^{2}\,,$ (11)
where $K_{\alpha}(x)$ is the modified Bessel function of the second kind. For convenience, let us define $R\equiv n_{\bar{n}}/n_{n}$ that has no momentum dependence.
For comparison, the non-degenerate approach gives an average antineutron fraction $R^{\prime}=2c^{2}s^{2}$ by neglecting the oscillating part Buccella:1987pv . Between the degenerate and non-degenerate cases, the difference in the antineutron fraction $R/R^{\prime}\equiv n_{\bar{n}}/n^{\prime}_{\bar{n}}$ can reach $\mathcal{O}(10^{3})$ orders with $(\mu-m)/T\sim 10^{4}$ which is a huge enhancement $e^{(\mu-m)/T}$ due to degeneracy. As shown in Fig. 1, the neutron degeneracy parameter $(\mu-m)/T$ decreases with radius but increases with time according to the NSCool simulation Page:2004fy ; NSCool program .
Figure 1: The degeneracy parameter $(\mu-m)/T$ of neutron matter inside NS as function of radius and cooling time.
Neutron Burning and Neutron Star Heating – A standing fraction of antineutron in the dense neutron environment as indicated by Eq. (10) would lead to constant burning of neutron matter by neutron-antineutron annihilation. In a single annihilation, the total nucleon number $N=N_{n}+N_{\bar{n}}$ decreases by two, $\mathrm{d}N=-2\langle\sigma v\rangle n_{n}n_{\bar{n}}\mathrm{d}tdV$, where $\langle\sigma v\rangle$ is the thermally averaged cross section $\sigma$ times the relative velocity $v$. Replacing the antineutron number density $n_{\bar{n}}=Rn_{n}$ gives,
$\frac{\mathrm{d}N}{\mathrm{d}t}=-\frac{2R}{1+R}n_{n}\braket{\sigma v}N\equiv-\Gamma N,$ (12)
where $\Gamma\equiv 2Rn_{n}\langle\sigma v\rangle/(1+R)$ takes the role of neutron burning rate.
Suppose the neutron burning does not consume a significant fraction of the NS mass, the neutron density $n_{n}$ and the annihilation rate $\Gamma$ can then be treated as constants if the temperature is also a constant. Then the total nucleon number decays as, $N(t)=N(0)e^{-\Gamma t}$. With $n_{n}=10^{38}~{}\rm{cm^{-3}}$ and $\braket{\sigma v}\approx 10^{-15}~{}\rm{cm^{3}/s}$ Buccella:1987pv , the neutron burning rate can reach $\Gamma\approx 2R\times 10^{23}~{}{\rm{s^{-1}}}$ where we have taken the approximation $R\ll 1$. For NS to exist for more than one million years, $\Gamma\lesssim 3\times 10^{-14}$ s-1, which would require $R\lesssim 10^{-37}$. The actual constraint is even much stronger since the temperature descreases very fast with time and the resulting degeneracy parameter $(\mu-m)/T$ can reach above $10^{5}\sim 10^{6}$ as shown in Fig. 1.
The annihilation products are mainly mesons such as $\pi^{0}$ and $\pi^{\pm}$. While the neutral $\pi^{0}$ finally decays into two photons to deposit all its energy as entropy in the NS, the charged $\pi^{\pm}$ produce both charged leptons (muon $\mu^{\pm}$ and electron/positron $e^{\mp}$) and neutrinos, $\pi^{\pm}\rightarrow\mu^{\pm}+\nu$ and $\mu^{\pm}\rightarrow e^{\pm}+2\nu$. The energies taken away by neutrinos are actually quite limited since the energetic pions from the neutron-antineutron annihilation first lose energy in dense material and then decay at rest. Similar thing happens for muons. So the neutrino energy spectrum from pion and muon decays are well defined. On average, neutrinos take way $\bar{E}_{\nu}=50.4$ MeV energy Liu:2021xzc . The other energies, $(m_{n}-\bar{E}_{\nu})\times dN/dt$, deposit to heat the NS and affect its cooling curve.
We implement the heating due to $n\bar{n}$ oscillation in NSCool Page:2004fy ; NSCool program to simulate the NS temperature evolution. As shown in Fig. 2, the heating effect by $n\bar{n}$ oscillation will prevent NS from cooling at some point. At the level of $10^{5}\sim 10^{6}$ years, the degeneracy parameter $(\mu-m)/T$ can reach $10^{5}\sim 10^{6}$. According to Eq. (11), the antineutron fraction $R_{0}$ is enhanced by a factor of $e^{2(\mu-m)/T}$ and hence its constraints can reach $\log_{10}R_{0}\sim-\mathcal{O}(10^{4})$ to be consistent with the observed NS cooling data Potekhin:2020ttj . When converted to the mixing angle, the constraint can reach $s^{2}\lesssim 10^{-\mathcal{O}(10^{4})}$.
Figure 2: The NS surface temperature data and cooling curves for both relativistic (solid) and non-relativistic (dashed) dispersion relations. Instead of the mixing angle $s^{2}$, the input parameter is taken as the antineutron density ratio $R_{0}$ with $T_{0}=9.9\times 10^{9}$ K at the beginning of the NS evolution and $n_{n}=5.8\times 10^{38}$ cm-3 around the NS center.
The mixing angle can be expressed approximately as $s\sim\delta m/\Delta H$ where $\Delta H\equiv H_{22}-H_{11}$ is the energy difference between neutron and antineutron in the range from 1 MeV Buccella:1987pv ; McKeen:2018xwc to 1 GeV Berezhiani:2018zvs . On the other hand, in some GUT models the cross mass term $\delta m$ is characterized by an effective mass scale $M_{X}$, $\delta m\sim\Lambda_{QCD}^{6}/M_{X}^{5}$ where $\Lambda_{\mathrm{QCD}}\simeq 180$ MeV Phillips:2014fgb . The enhanced constraint from NS heating can reach $M_{X}>10^{\mathcal{O}(1000)}$ GeV which is far beyond the Planck scale $\sim 10^{19}$ GeV.
Summary and Conclusions – Although the fermionic oscillation has been well established and experimental verified with neutrino oscillations, the previous study is only about the oscillation in vauum or with an effective matter potential at most. The description of $n\bar{n}$ oscillation is simply borrowed from its neutrino counterpart. However, most neutrons are confined inside nuclei or the compact NS with high degeneracy. It is then necessary for $n\bar{n}$ oscillation to consider the degenerate environment. Although our analysis takes the $n\bar{n}$ oscillation inside NS as illustration, the same formalism should also apply for nuclei.
Our result shows that this degeneracy effect can significantly change the neutron oscillation behavior. Especially, the probed GUT scale can go far beyond the Planck scale which means that those GUT models with $n\bar{n}$ oscillation are essentially excluded by the stability of NS. Otherwise, NS may explode when the degeneracy parameter $(\mu-m)/T$ significantly increases with much lower internal temperature towards the end of NS cooling curve. Similar conclusion should also apply for the neutron-mirror-neutron oscillation scenarios with cooling effect.
Acknowledgements
The authors would like to thank Bo-Wen Ouyang, Pedro Pasquini, Jiaming Zheng, and Albert Zhou for useful discussions. The authors are supported by the National Natural Science Foundation of China (12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604). SFG is also an affiliate member of Kavli IPMU, University of Tokyo. XYF, ZYG, and QHW are supported by the Zhiyuan Future Scholar program (ZIRC2024-10) of the SJTU Zhiyuan College.

### v2

@@
-Degeneracy Enhancement of Neutron-Antineutron Oscillation in Neutron Star
+Degeneracy Enhancement of Neutron Oscillation in Neutron Star
 Abstract
@@
 The subsequent neutron star heating can put an extremely stringent bound on the baryon number violating cross mass term between neutron and antineutron.
+The neutron-mirror-neutron oscillation in neutron star with cooling effect should have similar results.
 Introduction – The Grand Unification Theory (GUT) Georgi:1974sy ; Pati:1974yy ; Mohapatra:1974hk ; Fritzsch:1974nn is a fundamental theory that seeks to unify the electromagnetic, weak, and strong forces Croon:2019kpe .
@@
 This enhancement of $n\bar{n}$ oscillation is very significant and allows NS to provide much stronger constraint than previously expected.
-Fermion Oscillation with Degeneracy – The formalism for fermion oscillation in vacuum is well established for neutrinos Pontecorvo:1957cp ; MNS ; P67 .
+Fermion Oscillation with Degeneracy – The formalism for fermion oscillation in vacuum is well established Pontecorvo:1957cp ; MNS ; P67 .
 The initial flavor state $|\alpha\rangle$ can be decomposed into a linear combination of mass eigenstates $|i\rangle$, $|\alpha\rangle=\sum_{i}U^{*}_{\alpha i}|i\rangle$ with $U_{\alpha i}$ being the mixing matrix element Kayser:2004wmw .
@@
 Correspondingly, the original flavor state $|\alpha(t,L)\rangle$ also evolves with time and distance.
-For measurement also in a flavor state $|\beta\rangle$, the amplitude $A_{\beta\alpha}\equiv\langle\beta|\alpha(t,L)\rangle$ is no longer simply $\langle\beta|\alpha\rangle=\delta_{\alpha\beta}$ but
-$A_{\beta\alpha}=\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}.$ (1)
-With different evolution phases $\phi_{i}\neq\phi_{j}$, their interference leads to oscillatory behavior in the transition probability $P_{\alpha\beta}\equiv|A_{\beta\alpha}|^{2}$.
-We have kept our notations general in order to cover neutrino and other fermion oscillations.
-For neutrino oscillation, the individual evolution phase is $\phi_{i}=m^{2}_{i}L/2E_{\nu}$ and the oscillation phase is $\Delta\phi_{ij}=\Delta m^{2}_{ij}L/2E_{\nu}$ with $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$.
-One may see that the neutrino oscillation phenomena is actually a quantum interference over macroscopic distance.
+For measurement also in a flavor state $|\beta\rangle$, the transition probability $P_{\alpha\beta}$ is no longer simply $\delta_{\alpha\beta}$ but
+$P_{\beta\alpha}=\left|\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}\right|^{2}.$ (1)
+With different evolution phases $\phi_{i}\neq\phi_{j}$, their interference leads to oscillatory behavior in the transition probability $P_{\alpha\beta}$.
+For neutrino oscillation, the individual evolution phase is $\phi_{i}\equiv m^{2}_{i}L/2E_{\nu}$ and the oscillation phase is $\Delta\phi_{ij}=\Delta m^{2}_{ij}L/2E_{\nu}$ with $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$.
+Note that our notations in Eq.
+(1) are kept general to cover both neutrino and other fermion oscillations.
 The vacuum fermion oscillation is established by considering fermion propagation on the vacuum state.
@@
 In a degenerate environment, the vacuum state $|0\rangle$ should be replaced by the degenerate environment state $|\Omega\rangle$.
-In a concrete process, the fermion in propagation is created by a creation operator $a^{\dagger}_{\bm{p}i}$ and then propagate in the dense environment before being annihilated by $a_{\bm{p}i}$.
+The fermion in propagation is created by a creation operator $a^{\dagger}_{\bm{p}i}$ and then propagate in the dense environment before being annihilated by $a_{\bm{p}i}$.
 The corresponding amplitude takes the form as $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle$.
@@
 When applying to the environmental state $|\Omega\rangle$, the number density operator $\hat{n}_{\bm{p}i}$ should produce the particle momentum density, $\hat{n}_{\bm{p}i}|\Omega\rangle=n_{\bm{p}i}|\Omega\rangle$ with $n_{\bm{p}i}\equiv\int\mathrm{d}^{3}\bm{x}f_{i}(\bm{x},\bm{p})$ where $f_{i}(\bm{x},\bm{p})$ is the corresponding phase space distribution function.
-For a fermion, the creation and annihilation operators anticommute as, $\{a_{\bm{p}_{1}i},a_{\bm{p}_{2}i}^{\dagger}\}=(2\pi)^{3}\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$.
-Consequently, the product $a_{\bm{p}i}a^{\dagger}_{\bm{p}i}$ gives the Pauli blocking factor,
+For a fermion, the creation and annihilation operators anticommute as, $\{a_{\bm{p}_{1}i},a_{\bm{p}_{2}i}^{\dagger}\}=(2\pi)^{3}\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$  [^1] During revision, we notice the same $\delta^{(3)}$ function has also been obtained in Ghosh:2022nzo ; Barbosa:2024tty ..
+Then the product $a_{\bm{p}i}a^{\dagger}_{\bm{p}i}$ gives the Pauli blocking factor Barbosa:2024tty ,
 $\displaystyle{a_{\bm{p}i}a_{\bm{p}i}^{\dagger}|\Omega\rangle=\int\mathrm{d}^{3}\bm{x}[1-f_{i}(\bm{x},\bm{p})]}|\Omega\rangle,$ (2)
-where the $\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$ has been removed by momentum integration in the second quantized field operator, $\nu(x)\sim\int\mathrm{d}^{4}p(aue^{-ip\cdot x}+b^{\dagger}ve^{ip\cdot x})$.
-Taking the neutrino oscillation as a concrete example, the whole scattering amplitude of production, propagation, and detection takes the form as,
-$\mathcal{M}$ $=\langle f,\Omega|(\bar{\ell}_{\beta}\gamma^{\mu}P_{L}W_{\mu}^{-}\nu_{\beta})(\bar{\nu}_{\alpha}\gamma^{\nu}P_{L}\ell_{\alpha}W_{\nu}^{+})|i,\Omega\rangle$ (3a)
-$\equiv\mathcal{M}_{d}\left[\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle\right]\mathcal{M}_{p},$ (3b)
-where $W^{\pm}$ is the $W$ gauge boson and $\ell_{\alpha}$ the charged lepton of flavor $\alpha$.
-For convenience, the weak coupling constant has been omitted.
-The two neutrino spinors $\bar{u}$ and $u$ have been absorpted into the production and detection matrix elements $\mathcal{M}_{p}$ and $\mathcal{M}_{d}$, respectively.
-In between, the creation and annihilation operators sandwiched by the thermal environment state $|\Omega\rangle$ would give the Pauli exclusion factor, $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle=1-f_{i}$.
-Although illustrated with neutrino oscillation, this feature should apply generally to any fermion oscillation.
-The oscillation amplitude in Eq. (1) then generalizes to
-$A_{\beta\alpha}=\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}(1-f_{i}).$ (4)
-The Pauli suppression factor $1-f_{i}$ represents the degeneracy effect brought by the environment.
-With an empty environment as vacuum ($f_{i}=0$), the amplitude reduces to the vacuum oscillation case as shown in Eq.
+where the $\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$ has been removed by momentum integration in the second quantized fermion field operator, $\psi(x)\sim\int\mathrm{d}^{4}p(aue^{-ip\cdot x}+b^{\dagger}ve^{ip\cdot x})$.
+The scattering amplitude for the whole process composed of production, propagation, and detection takes the form as,
+$\mathcal{M}_{\beta\alpha}$ $\equiv\mathcal{M}_{d}\left[\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle\right]\mathcal{M}_{p},$ (3)
+The fermion oscillation process is a very special type with the intermediate particle being on shell and the particle momentum does not change.
+These intermediate particles remain in coherence and evolve according to definite dispersion relation.
+The creation and annihilation operators from the intermediate fields applies on the thermal vacuum $|\Omega\rangle$ to produce the Pauli exclusion factor, $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle=1-f_{i}$.
+The two on-shell spinors $\bar{u}$ and $u$ have been absorpted into the production and detection matrix elements $\mathcal{M}_{p}$ and $\mathcal{M}_{d}$, respectively.
+Attaching a Pauli blocking factor $1-f_{i}$ with an on-shell intermediate particle is fully consistent with the thermal quantum field theory (QFT) Lundberg:2020mwu .
+The fermion propagator in an ensemble of particles, $S_{F}(p)\propto i/(p^{2}-m^{2}+i\epsilon)-2\pi\delta(p^{2}-m^{2})\Theta(p^{0})f(\bm{p})$ Ghosh:2022nzo ; Notzold:1987ik would reduce to
+$S_{F}$ $\propto\int\frac{d^{3}\bm{p}}{(2\pi)^{3}}\frac{1}{2E_{\bm{p}}}e^{-ip\cdot(x-y)}[1-f(\bm{p})]\left(\not{p}+m_{\nu}\right),$ (4)
+after Fourier transformation $\int d^{4}p/(2\pi)^{4}$.
+The first term in the square bracket comes from the ordinary propagator $i/(p^{2}-m^{2}+i\epsilon)$ by extracting the physical pole while the $f(\bm{p})$ term from the thermal contribution after integrating away the on-shell $\delta(p^{2}-m^{2})$ function.
+Treating the $1-f$ factor as field strength renormalization, our result can be readily obtained with the quantum kinetics theory Sigl:1993ctk ; Volpe:2023met .
+From the mass matrix diagonalization, the flavor fields $\psi_{\alpha}$ connect with the mass eigen-fields $\psi_{i}$ via a unitarity mixing matrix $U$, $\psi_{\alpha}=\sum_{i}U_{\alpha i}\psi_{i}$.
+By field redefinition, $\psi_{i}=\sqrt{Z_{i}}\psi^{\prime}_{i}$ with $Z_{i}\equiv 1-f_{i}$, the canonical field $\psi^{\prime}_{i}$ develops evolution phase $e^{i\phi_{i}(t,L)}$.
+Then the scattering matrix element should contain a factor of $U_{\beta i}(1-f_{i})e^{i\phi_{i}}U^{*}_{\alpha i}$ in the flavor basis, which is consistent with Eq.
 (3).
-On the other hand, a fully degenerate Fermi gas $f_{i}=1$ does not allow oscillation to occupy the corresponding mass eigenstate.
-With Pauli suppression factor, the amplitude in Eq.
-(4) can no longer reflect unitarity.
-Sticking to the usual definition of oscillation probability, $P_{\alpha\beta}\equiv|A_{\beta\alpha}|^{2}$, the total probability, $P_{\text{tot},\alpha}\equiv\sum_{\gamma}P_{\alpha\gamma}\neq 1$, is no longer one.
-It seems that the oscillating fermion would somehow disappear which is contradictory to the intuitive physical picture.
-However, oscillation in a degenerate environment does not introduce any mechanism for a particle to disappear.
-So unitarity has to be maintained.
-To resolve this contradiction, one needs to resort to the more intuitive definition of oscillation probability.
-The original definition of probability or chance is actually the fraction of events.
-To be more exact, the oscillation probability $P_{\alpha\beta}$ should be the ratio between the number of $\beta$ flavor particles that can be detected after oscillation and the original number of $\alpha$ flavor ones without oscillation.
-It is then necessary to normalize the oscillation probability as $\mathcal{P}_{\alpha\beta}\equiv|A_{\alpha\beta}|^{2}/P_{\text{tot}}$ by hand.
-Note that the amplitude $A_{\alpha\beta}$ in Eq.
-(4) does not reduce to $\delta_{\alpha\beta}$ in the zero distance limit since the Pauli blocking is already in effect even if the oscillation does not really happen.
-In a degenerate environment, an $\alpha$ flavor state from production cannot contain the existing mass eigenstate.
+In the QFT language, the detected particle number of flavor $\beta$ is proportional to $N_{\beta}\propto|\mathcal{M}_{\beta\alpha}|^{2}$.
+To be more specific, the total particle number without propagation ($x=y$) is proportional to $N\propto\sum_{\beta}|\mathcal{M}_{\beta\alpha}(x=y)|^{2}$.
+After propagating over a nonzero space-time interval $x-y$, the detected particle number of flavor $\beta$ is proportional to $N_{\beta}(x-y)\propto|\mathcal{M}_{\beta\alpha}(x-y)|^{2}$.
+In other words, a fraction of $N_{\beta}(x-y)/N$ has oscillated to flavor $\beta$ which gives the probability for a neutron to be probed as an antineutron,
+$P_{\alpha\beta}\equiv\frac{\left|\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}(1-f_{i})\right|^{2}}{\sum_{\beta}\left|\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}(1-f_{i})\right|^{2}}.$ (5)
+With an empty environment ($f_{i}=0$), the transition probability reduces to the vacuum case as shown in Eq.
+(1).
+The same formalism of degenerate oscillation should also apply for bosons with the Pauli blocking factor $1-f_{i}$ replaced by the Bose enhancement factor $1+f_{i}$.
 Neutron-Antineutron Oscillation with Degeneracy – Being neutral particles, neutron and antineutron can have cross mass term Phillips:2014fgb .
@@
 The diagonal elements $H_{11}$ and $H_{22}$ are the neutron and antineutron energies Phillips:2014fgb .
-The cross mass $\delta m$ will induce the $n$-$\bar{n}$ mixing, $\tan 2\theta=2\delta m/(H_{22}-H_{11})$,
-$U=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix},\quad\begin{pmatrix}n_{1}\\ n_{2}\end{pmatrix}=U\begin{pmatrix}n\\ \bar{n}\end{pmatrix},$ (6)
+The Majorana mass $\delta m$ that violates the baryon number by 2 will induce the $n$-$\bar{n}$ mixing, $\tan 2\theta=2\delta m/(H_{22}-H_{11})$,
+$\begin{pmatrix}n_{1}\\ n_{2}\end{pmatrix}=U\begin{pmatrix}n\\ \bar{n}\end{pmatrix},\quad U=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix},$ (7)
 where $n_{1}$ and $n_{2}$ are the two mass eigen-fields.
@@
 The energy difference between the mass eigenstates is simply denoted as $\Delta E\equiv\sqrt{(H_{11}-H_{22})^{2}+4\delta m^{2}}$ Phillips:2014fgb .
-So the $n\bar{n}$ oscillation probabilities are
-$P_{n\bar{n}}$ $=c^{2}s^{2}(1-f_{1})^{2}+s^{2}c^{2}(1-f_{2})^{2}$
-$-2c^{2}s^{2}(1-f_{1})(1-f_{2})\cos(\Delta Et),$ (7a)
-$P_{nn}$ $=c^{4}(1-f_{1})^{2}+s^{4}(1-f_{2})^{2}$
-$+2c^{2}s^{2}(1-f_{1})(1-f_{2})\cos(\Delta Et),$ (7b)
+The oscillation probability takes the form as,
+$P_{n\bar{n}}$ $=\frac{c^{2}s^{2}(1-f_{1})^{2}+c^{2}s^{2}(1-f_{2})^{2}}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}}$
+$-\frac{2(1-f_{1})(1-f_{2})c^{2}s^{2}\cos(\Delta Et)}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}},$ (8)
 where $f_{1}$ and $f_{2}$ are the phase space distribution functions of the two mass eigenstates.
-Note that the time dependence only appears in the $\cos(\Delta Et)$ term and the norminal total probability $P_{\rm tot}\equiv P_{nn}+P_{n\bar{n}}$ is indeed time-independent which is a necessary feature for our argument of probability normalization/definition above.
-The two probabilities reduce to the non-degenerate case with both $f_{1}$ and $f_{2}$ approaching 00.
-The normalized oscillation probability should take the form as,
-$\mathcal{P}_{n\bar{n}}$ $=\frac{c^{2}s^{2}(1-f_{1})^{2}+c^{2}s^{2}(1-f_{2})^{2}}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}}$
-$-\frac{2(1-f_{1})(1-f_{2})c^{2}s^{2}\cos(\Delta Et)}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}},$ (8)
-with $\mathcal{P}_{nn}\equiv P_{nn}/P_{\rm tot}$ being omitted for simplicity but can be readily reconstructed.
-The total probability $P_{\rm tot}$ in the denominator reduces to 1 in the non-degenerate limit.
-It is interesting to see that the phase space distribution function appears in the denominator as $1-f_{i}$.
+Note that the phase space distribution function appears in the denominator as $1-f_{i}$.
 In a dense environment with degeneracy, the phase space distribution function $f_{i}$ can approach 1 below the Fermi surface.
@@
 Consequently, both $1-f_{1}$ and $f_{2}$ are negligibly small.
-Between the two tiny terms $s$ and $1-f_{1}$ in the denominator, the case $s\gg 1-f_{1}$ would lead to significant transition rate $\mathcal{P}_{n\bar{n}}\approx c^{2}$ which means degenerate neutrons would instantly transit to antineutrons.
+Between the two tiny terms $s$ and $1-f_{1}$ in the denominator, the case $s\gg 1-f_{1}$ would lead to a large $P_{n\bar{n}}\approx c^{2}$ which means degenerate neutrons would instantly transit to antineutrons.
 The situation $s\sim 1-f_{1}$ has also the same conclusion.
@@
 This already indicates that the degenerate $n\bar{n}$ oscillation in NS can give a very much enhanced constraint.
-To be more exact, in the limit of $s\ll 1-f_{1}\ll 1$ and $f_{2}\ll 1$, the transition rate from neutron to antineutron becomes
-$\mathcal{P}_{n\bar{n}}\approx\frac{s^{2}}{(1-f_{1})^{2}}\equiv s^{2}_{\rm eff},\qquad\mathcal{P}_{nn}\approx c^{2}.$ (9)
-It is interesting to observe that the transition rate $\mathcal{P}_{n\bar{n}}$ is enhanced by degeneracy.
-This is consistent with the intuitive picture that it is much easier for neutrons in a degenerate neutron environment to transit to other particles without Pauli blocking.
-The effective mixing $s^{2}_{\rm eff}$ should be much larger than the vacuum counterpart $s^{2}$.
-Switching neutron with antineutron and hence $f_{1}$ with $f_{2}$, the reversed transition probability $\mathcal{P}_{\bar{n}n}\approx s^{2}$ is not enhanced.
-The bidirectional oscillation has imbalanced rates to derive the overall transition from the degenerate neutrons to the non-degenerate antineutrons.
-Standing Fraction of Antineutron in Neutron Star – Another important feature is that the time dependence of the $n\bar{n}$ oscillation probability $\mathcal{P}_{n\bar{n}}$, which is $2c^{2}s^{2}\sin^{2}(\Delta Et/2)$ for the vacuum case, is now suppressed by extra $1-f_{1}$.
+Figure 1: Illustration of the degenerate $n$-$\bar{n}$ oscillation.
+In the limit of $s\ll 1-f_{1}\ll 1$ and $f_{2}\ll 1$, the transition rate from neutron to antineutron becomes
+$P_{n\bar{n}}\approx\frac{s^{2}}{(1-f_{1})^{2}}\equiv s^{2}_{\rm eff},\qquad P_{nn}\approx c^{2}.$ (9)
+As shown in Fig.
+1, although the neutron state ($n$) is mainly composed of $n_{1}$, its coefficient $\cos\theta$ receives extra suppression $1-f_{1}$ due to Pauli blocking.
+So the relative fraction of $n_{2}$ and subsequently the chance for the antineutron ($\bar{n}$) to appear is enhanced.
+It is much easier for neutrons in a degenerate neutron environment to transit to antineutrons without Pauli blocking.
+Switching neutron with antineutron and hence $f_{1}$ with $f_{2}$, the reversed transition probability $P_{\bar{n}n}\approx s^{2}$ is not enhanced.
+The bidirectional oscillation has imbalanced rates to drive the overall transition from the degenerate neutrons to the non-degenerate antineutrons.
+Standing Fraction of Antineutron in Neutron Star – Another important feature is that the time dependence of the $n\bar{n}$ oscillation probability $P_{n\bar{n}}$, which is $2c^{2}s^{2}\sin^{2}(\Delta Et/2)$ for the vacuum case, is now suppressed by extra $1-f_{1}$.
 In other words, the transition happens instantaneously.
-In an extremely degenerate environment such as NS, neutrons will transit into antineutrons so rapidly that the ratio between their distribution functions $f_{\bar{n}}/f_{n}$ stays the same,
-$R(\bm{p})\equiv\frac{f_{\bar{n}}}{f_{n}}=\frac{\mathcal{P}_{n\bar{n}}}{\mathcal{P}_{nn}}\approx\frac{\tan^{2}\theta}{[1-f_{1}(\bm{p})]^{2}},$ (10)
+Further more, the environment is full of neutrons such that no particle can freely propagate for sizable distance.
+It is unavoidable to consider the zero time/distance limit.
+In an extremely degenerate environment, neutrons transit into antineutrons so rapidly that the ratio between their distribution functions $f_{\bar{n}}/f_{n}$ stays the same,
+$R(\bm{p})\equiv\frac{f_{\bar{n}}}{f_{n}}=\frac{P_{n\bar{n}}}{P_{nn}}\approx\frac{\tan^{2}\theta}{[1-f_{1}(\bm{p})]^{2}},$ (10)
 which actually reduces to the effective mixing $s^{2}_{\rm eff}$.
@@
 Even directly taking the zero time limit of Eq.
-(8), $\mathcal{P}_{n\bar{n}}(t=0)\approx s^{2}(f_{1}-f_{2})^{2}/(1-f_{1})^{2}$ where the approximation only applies for the denominator, Eq.
+(8), $P_{n\bar{n}}(t=0)\approx s^{2}(f_{1}-f_{2})^{2}/(1-f_{1})^{2}$ where the approximation only applies for the denominator, Eq.
 (10) can be readily reproduced with $(f_{1}-f_{2})^{2}\approx 1$.
 In other words, the production process is already affected by the degenerate environment as pointed out earlier.
-The neutron transition to antineutron can instantly happen and remain the same antienutron fraction over time.
+The neutron transition to antineutron can instantly happen and remain the same antineutron fraction over time.
 Being in thermal equilibrium as maintained inside NS, neutrons follow the Fermi-Dirac distribution, $f_{1}(\bm{p})=1/[e^{(\varepsilon_{n}(\bm{p})-\mu)/T}+1]$ where $\varepsilon_{n}(\bm{p})=\sqrt{\bm{p}^{2}+m^{2}}$ is the neutron energy with momentum $\bm{p}$ and mass $m$.
-With a very low internal temperature $T\sim 7$ keV at the age of $10^{5}$ years Potekhin:2020ttj , the corresponding neutron Fermi momentum $k_{F}=(3\pi^{2}n_{n})^{1/3}$ and chemical potential $\mu_{n}=\sqrt{k_{F}^{2}+m^{2}}$ can be readily derived from the neutron density $n_{n}$ assuming full degeneracy.
+With a very low internal temperature $T\sim 7$ keV at the age of $10^{5}$ years Potekhin:2020ttj , the corresponding neutron Fermi momentum $k_{F}=(3\pi^{2}n_{n})^{1/3}$ and chemical potential $\mu=\sqrt{k_{F}^{2}+m^{2}}$ can be readily derived from the neutron density $n_{n}$ assuming full degeneracy.
 We consider a typical NS with mass $M=1.4M_{\odot}$ (solar mass $M_{\odot}$) and radius $R_{n}=10$ km Lattimer:2000nx .
@@
 It is fully consistent to say that the distribution function is very close to 1 below the Fermi surface, $1-f_{1}(\bm{p})\approx e^{[\varepsilon_{n}(\bm{p})-\mu]/T}\ll 1$.
+Although neutrons can form superfluid, it can only happen around the Fermi surface with a gap $\sim 1$ MeV Haskell:2017lkl much smaller than the chemical potential $\mu-m$ and Fermi momentum.
+In addition, $f_{1}$ deviates from 1 in this region and hence the $n\bar{n}$ oscillation would not experience degeneracy enhancement.
 Since the first mass eigenstate is mainly neutron, $n_{1}=c\,n+s\,\bar{n}$ according to Eq.
@@
 As shown in Fig.
-1, the neutron degeneracy parameter $(\mu-m)/T$ decreases with radius but increases with time according to the NSCool simulation Page:2004fy ; NSCool program .
-Figure 1: The degeneracy parameter $(\mu-m)/T$ of neutron matter inside NS as function of radius and cooling time.
+2, the neutron degeneracy parameter $(\mu-m)/T$ decreases with radius but increases with time according to NSCool Page:2004fy ; NSCool program .
+Figure 2: The degeneracy parameter $(\mu-m)/T$ of neutron matter inside NS without considering $n\bar{n}$ annihilation
+as function of radius and cooling time.
 Neutron Burning and Neutron Star Heating – A standing fraction of antineutron in the dense neutron environment as indicated by Eq.
 (10) would lead to constant burning of neutron matter by neutron-antineutron annihilation.
-In a single annihilation, the total nucleon number $N=N_{n}+N_{\bar{n}}$ decreases by two, $\mathrm{d}N=-2\langle\sigma v\rangle n_{n}n_{\bar{n}}\mathrm{d}tdV$, where $\langle\sigma v\rangle$ is the thermally averaged cross section $\sigma$ times the relative velocity $v$.
+In a single annihilation, the total nucleon number $N=N_{n}+N_{\bar{n}}$ decreases by two, $\mathrm{d}N=-2\langle\sigma v\rangle n_{n}n_{\bar{n}}\mathrm{d}t\mathrm{d}V$, where $\langle\sigma v\rangle$ is the thermally averaged cross section $\sigma$ times the relative velocity $v$.
 Replacing the antineutron number density $n_{\bar{n}}=Rn_{n}$ gives,
 $\frac{\mathrm{d}N}{\mathrm{d}t}=-\frac{2R}{1+R}n_{n}\braket{\sigma v}N\equiv-\Gamma N,$ (12)
-where $\Gamma\equiv 2Rn_{n}\langle\sigma v\rangle/(1+R)$ takes the role of neutron burning rate.
+where $\Gamma\equiv 2Rn_{n}\langle\sigma v\rangle/(1+R)$ is the neutron burning rate.
 Suppose the neutron burning does not consume a significant fraction of the NS mass, the neutron density $n_{n}$ and the annihilation rate $\Gamma$ can then be treated as constants if the temperature is also a constant.
 Then the total nucleon number decays as, $N(t)=N(0)e^{-\Gamma t}$.
-With $n_{n}=10^{38}~{}\rm{cm^{-3}}$ and $\braket{\sigma v}\approx 10^{-15}~{}\rm{cm^{3}/s}$ Buccella:1987pv , the neutron burning rate can reach $\Gamma\approx 2R\times 10^{23}~{}{\rm{s^{-1}}}$ where we have taken the approximation $R\ll 1$.
-For NS to exist for more than one million years, $\Gamma\lesssim 3\times 10^{-14}$ s-1, which would require $R\lesssim 10^{-37}$.
-The actual constraint is even much stronger since the temperature descreases very fast with time and the resulting degeneracy parameter $(\mu-m)/T$ can reach above $10^{5}\sim 10^{6}$ as shown in Fig.
-1.
+With $n_{n}=10^{38}\,\rm{cm^{-3}}$ and $\braket{\sigma v}\approx 10^{-15}\,\rm{cm^{3}/s}$ Buccella:1987pv , the neutron burning rate can reach $\Gamma\approx 2R\times 10^{23}\,{\rm{s^{-1}}}$ where we have taken the approximation $R\ll 1$.
+For NS to exist for more than one million years, $\Gamma\lesssim 3\times 10^{-14}$ s-1, which would require a lose constraint $R\lesssim 10^{-37}$.
+The actual constraint is even much stronger since the temperature descreases very fast with time and the resulting degeneracy parameter $(\mu-m)/T$ can reach above $10^{5}\sim 10^{6}$ Bertoni:2013bsa as shown in Fig.
+2.
 The annihilation products are mainly mesons such as $\pi^{0}$ and $\pi^{\pm}$.
@@
 The other energies, $(m_{n}-\bar{E}_{\nu})\times dN/dt$, deposit to heat the NS and affect its cooling curve.
+Figure 3: The red-shifted NS surface temperature data and cooling curves for both relativistic (solid) and non-relativistic (dashed) dispersion relations.
+Instead of the mixing angle $s^{2}$, the input parameter is taken as the antineutron density ratio $R_{0}\equiv R(n_{0},T_{0})$ with $T_{0}=9.9\times 10^{9}$ K at the beginning of the NS evolution and $n_{0}=5.8\times 10^{38}$ cm-3 around the NS center.
 We implement the heating due to $n\bar{n}$ oscillation in NSCool Page:2004fy ; NSCool program to simulate the NS temperature evolution.
 As shown in Fig.
-2, the heating effect by $n\bar{n}$ oscillation will prevent NS from cooling at some point.
-At the level of $10^{5}\sim 10^{6}$ years, the degeneracy parameter $(\mu-m)/T$ can reach $10^{5}\sim 10^{6}$.
-According to Eq.
-(11), the antineutron fraction $R_{0}$ is enhanced by a factor of $e^{2(\mu-m)/T}$ and hence its constraints can reach $\log_{10}R_{0}\sim-\mathcal{O}(10^{4})$ to be consistent with the observed NS cooling data Potekhin:2020ttj .
+3, the heating effect by $n\bar{n}$ oscillation will prevent NS from cooling at some point such that the cooling curve becomes flat.
+To be consistent with the observed data points Potekhin:2020ttj , the initial parameter $R_{0}$ should not exceed $\log_{10}R_{0}\sim-\mathcal{O}(10^{4})$.
+With $(\mu-m)/T$ reaching $10^{6}$ Bertoni:2013bsa around $10^{5}$ years, the physical antineutron fraction $R\approx(T/T_{0})^{3/2}e^{2(\mu-m)/T-2(\mu_{0}-m)/T_{0}}R_{0}$ is enhanced to $R\lesssim 10^{-43}$.
+Correspondingly, the heat release $2Rn^{2}_{n}\langle\sigma v\rangle Vm$ in the NS core with a radius of 2 km can reach $10^{24}$ W to far exceed the cooling power $4\pi R^{2}\sigma T^{4}\sim 10^{21}$ W at the NS surface.
 When converted to the mixing angle, the constraint can reach $s^{2}\lesssim 10^{-\mathcal{O}(10^{4})}$.
-Figure 2: The NS surface temperature data and cooling curves for both relativistic (solid) and non-relativistic (dashed) dispersion relations.
-Instead of the mixing angle $s^{2}$, the input parameter is taken as the antineutron density ratio $R_{0}$ with $T_{0}=9.9\times 10^{9}$ K at the beginning of the NS evolution and $n_{n}=5.8\times 10^{38}$ cm-3 around the NS center.
 The mixing angle can be expressed approximately as $s\sim\delta m/\Delta H$ where $\Delta H\equiv H_{22}-H_{11}$ is the energy difference between neutron and antineutron in the range from 1 MeV Buccella:1987pv ; McKeen:2018xwc to 1 GeV Berezhiani:2018zvs .
@@
 The enhanced constraint from NS heating can reach $M_{X}>10^{\mathcal{O}(1000)}$ GeV which is far beyond the Planck scale $\sim 10^{19}$ GeV.
-Summary and Conclusions – Although the fermionic oscillation has been well established and experimental verified with neutrino oscillations, the previous study is only about the oscillation in vauum or with an effective matter potential at most.
-The description of $n\bar{n}$ oscillation is simply borrowed from its neutrino counterpart.
-However, most neutrons are confined inside nuclei or the compact NS with high degeneracy.
-It is then necessary for $n\bar{n}$ oscillation to consider the degenerate environment.
-Although our analysis takes the $n\bar{n}$ oscillation inside NS as illustration, the same formalism should also apply for nuclei.
+Summary and Conclusions – We establish the formalism of degenerate oscillation in the QFT language with three different approaches of second quantization, thermal QFT, and kinetics theory.
+Although it can also apply for the bosonic case, we take the neutron-antineutron oscillation in NS for a concrete illustration.
 Our result shows that this degeneracy effect can significantly change the neutron oscillation behavior.
@@
 Acknowledgements
-The authors would like to thank Bo-Wen Ouyang, Pedro Pasquini, Jiaming Zheng, and Albert Zhou for useful discussions.
+The authors would like to thank Bo-Wen Ouyang, Pedro Pasquini, Liang Tan, Jiaming Zheng, Albert Zhou, and Wei Zuo for useful discussions.
+SFG is particularly grateful to Luke Johns, Georg Raffelt, Guenter Sigl, and Maria Cristina Volpe for useful comments and discussions.
 The authors are supported by the National Natural Science Foundation of China (12375101, 12090060 and 12090064) and the SJTU Double First Class start-up fund (WF220442604).